# AIGCodeSet: A New Annotated Dataset for AI Generated Code Detection

### v1

@@ Abstract
With the rapid advancement of LLM models, they have become widely useful in various fields. While these AI systems can be used for code generation, significantly simplifying and accelerating the tasks of developers, their use for students to do assignments has raised ethical questions in the field of education. In this context, determining the author of a particular code becomes important. In this study, we introduce AIGCodeSet, a dataset for AI-generated code detection tasks, specifically for the Python programming language. We obtain the problem descriptions and human-written codes from the CodeNet dataset. Using the problem descriptions, we generate AI-written codes with CodeLlama 34B, Codestral 22B, and Gemini 1.5 Flash models in three approaches: i) generating code from the problem description alone, ii) generating code using the description along with human-written source code containing runtime errors, and iii) generating code using the problem description and human-written code that resulted in wrong answers. Lastly, we conducted a post-processing step to eliminate LLM output irrelevant to code snippets. Overall, AIGCodeSet consists of 2,828 AI-generated and 4,755 human-written code snippets. We share our code with the research community to support studies on this important topic and provide performance results for baseline AI-generated code detection methods.
@@ 1 Introduction
With the impressive advancements in generative artificial intelligence (AI) technologies, large language models (LLM) become one of the main resources for coding. AI can enable software developers to complete coding tasks up to twice as fast, according to a recent study  [^1] https://www.mckinsey.com/capabilities/mckinsey-digital/our-insights/unleashing-developer-productivity-with-generative-ai. Similarly, a survey conducted by Stack Overflow [^2] https://stackoverflow.blog/2023/06/14/hype-or-not-developers-have-something-to-say-about-ai revealed that developers reported a 33% increase in productivity when using AI-powered tools. Furthermore, these tools can be beneficial for students as they offer a range of valuable resources, including sample solutions, various approaches to problem-solving, code reviews, exercise generation, code explanations, and illustrative examples Becker et al. (2023).
Despite the benefits of generative AI in programming, there are several concerns such as the risk of academic misconduct, plagiarism, issues with code reuse, and licensing. Furthermore, utilizing LLMs for coding poses significant risks as it may sometimes lack security, potentially introducing vulnerabilities Ambati et al. (2024). Perry et al. (2023) report that AI coding tools can generate insecure code in laboratory settings, raising considerable concerns about their use in real-world scenarios. Additionally, Jesse et al. (2023) highlights that AI-generated code using LLMs may contain errors, bugs, or inefficiencies due to the model’s lack of real-time testing and validation. Also, constantly relying on it can lead to the atrophy of coding skills Prather et al. (2023) or cause real developers to lose their jobs Barenkamp et al. (2020).
Due to the potential negative impact of these tools, several researchers focused on how to detect AI-generated code Hou et al. . The majority of these studies use LLMs developed by OpenAI and focus on using LLMs to generate code from scratch using the problem definition. However, LLMs can be used for other purposes in programming such as fixing an error in a given code snippet.
In this study, we introduce AIGCodeSet which is an annotated dataset for detecting AI-generated codes. We focus on Python due to its popularity. We first obtained 317 programming problems and corresponding submissions in Python from IBM’s CodeNet Puri et al. (2021). We sample submissions in three types: i) successful submissions, ii) submissions with runtime errors, and iii) submissions that produce incorrect output. Next, we use CodeLLAMA-34B, Codestral-22B, and Gemini 1.5 Flash LLMs to generate code separately. We target
We use LLMs in three different usage scenarios: i) generating a code from scratch using the problem definition, ii) fixing a given code with runtime error, and iii) correcting a given code with erroneous output. We use our sampled human-written codes to generate the codes for the last two scenarios. After generating the codes, we conduct a manual post-processing step to remove LLM output not related to the code (e.g, explainatiom about the code generation task).
Eventually, AIGCodeSet consists of 4,755 human-written python code snippets and 2,828 AI-generated ones, covering 317 programming tasks, three LLMs, and three LLM usage scenarios.
The contributions of our work are as follows.
  • We construct AIGCodeSet, an annotated dataset for AI generated code detection. AIGCodeSet covers CodeLLAMA-34B, Codestral-22B, and Gemini 1.5. To our knowledge, none of these models have been used in prior work [^3] Other variants of CodeLLAMA and Gemini have been used in a few studies. More importantly, AIGCodeSet cover utilization of LLMs to fix a given code but no prior work covers this usage scenario. We believe that AIGCodeSet will be a beneficial resource for the research community.
  • We provide performance results for baseline systems on our dataset and examine how their performance vary across LLMs and LLM usage scenarios.
The rest of the paper is structured as follows. We discuss related work in Section 2. In Section 3, we explain how we construct our dataset. We present and discuss our experimental results in Section 4 and conclude in Section 5.
@@ 2 Related Work
Despite relatively recent advances in large language models (LLMs) for code generation, this emerging area has garnered significant attention from researchers focusing on various aspects such as developing methods to detect AI-generated code Xu and Sheng (2024); Oedingen et al. (2024); Bukhari et al. (2023); Hoq et al. (2024); Idialu et al. (2024), evaluating existing AI-generated code detectors Pan et al. (2024); Wang et al. (2023), and detecting vulnerabilities in AI-generated codes Cotroneo et al. (2025); Wang et al. (2024). The studies differ in terms of the languages investigated, number of problems, size of datasets, LLMs employed, and source of the data.
The comparison of research findings is complicated due to the limited dataset resources comparing the diversity in LLMs, programming languages, and coding problems. There are few studies introducing datasets specifically for AI-generated code. Tihanyi et al. (2023) created a dataset containing 112,000 C language code samples generated using GPT-3.5 Turbo. Similarly, Wang et al. (2024) developed the CodeSecEval dataset, utilizing several LLMs, including multiple GPT models, CodeLlama-7B, and Claude 3 Opus, for Python code. Both datasets primarily address the security of generated code rather than the detection of AI-generated code.
Research on AI-generated code has mainly focused on specific programming languages. Python has been the most studied Pan et al. (2024); Cotroneo et al. (2025); Oedingen et al. (2024); Idialu et al. (2024); Wang et al. (2024), followed by C Bukhari et al. (2023); Tihanyi et al. (2023), and Java Hoq et al. (2024). Some studies have explored multiple languages; for instance, Xu and Sheng (2024) examined C, C++, C#, Java, JavaScript, and Python, while Wang et al. (2023) investigated a set including Ruby, JavaScript, Go, Python, Java, and PHP. In our study, we also focus on Python due to its popularity.
In creating a dataset for AI generated code detection, it is important to determine the problems to be covered and how to obtain the human-written codes. Existing studies use various resources for problem definitions and human-written codes. For instance, Pan et al. (2024) utilize codes and problems from Kaggle, Quescol, and LeetCode; Hoq et al. (2024) incorporate CodeWorkout; and Idialu et al. (2024) use CodeChef. Similar to our approach, Xu and Sheng (2024) utilize the CodeNet dataset, applying criteria such as selecting code with line lengths between 10 and 100, an alphanumeric character fraction greater than 0.25, and excluding all comments and duplicate files. In our study, we focus on problems which has sufficient successful and unsuccessful (i.e., runtime error or incorrect output) codes.
Regarding the LLMs used for AI-generated code studies, OpenAI’s models are the most popular ones such as ChatGPT Wang et al. (2023); Liu et al. (2024); Suh et al. (2024); Pan et al. (2024); Hoq et al. (2024), Davinci Xu and Sheng (2024); Bukhari et al. (2023), GPT-4 Idialu et al. (2024); Suh et al. (2024), and tools incorporating GPT-4, such as GitHub Co-Pilot and Microsoft Co-Pilot Cotroneo et al. (2025). Other LLMs include Google’s Bard Ambati et al. (2024), Gemini Suh et al. (2024); Cotroneo et al. (2025), CodeLlama-7B and Claude 3 Opus Wang et al. (2024).
Our study diverges from existing research in two significant aspects. First, to the best of our knowledge, no prior work has utilized Codestral or Code Llama 34B for detecting AI-generated code. Second, while most previous studies have focused on generating code from problem definitions, none have explored detecting AI-generated code specifically in scenarios where LLMs are employed to correct errors in provided human-written code for a programming problem.
@@
3 AIGCodeSet
In developing a dataset for detecting AI-generated code, the main requirement is to include both human-authored and AI-generated code samples. However, to ensure the dataset is of high quality and supports effective model training and reliable evaluation, several key considerations must be addressed. Specifically, it is essential to cover a wide range of coding problems and styles. Moreover, including both human-authored and AI-generated code for the same problem will allow for a more effective comparison of their differences. In addition, it is important to consider the level of involvement of AI tools and how they are used, as different individuals may use these tools in varying ways. Lastly, the dataset should cover various LLMs, as individuals may rely on different tools to solve their coding problems. Now, we explain our approach to constructing AIGCodeSet that aims to meet these objectives.
3.1 Acquiring Human Written Codes
Following prior work, we utilize IBM’s CodeNet dataset Puri et al. (2021) as a source for human-written code. This dataset contains 14 million code examples spanning approximately 4,000 coding problems across 55 programming languages. However, to align with our objectives and generate corresponding AI-generated versions at a reasonable cost, we selectively filter the dataset, retaining only the code samples most relevant to our study.
In our study, we focus specifically on the Python programming language due to its widespread popularity among programmers. The official webpage of the CodeNet dataset [^4] https://developer.ibm.com/exchanges/data/all/project-codenet/ also provides Python benchmark data that has 801 unique problems in Python. Upon examining the data, we observed that the problem descriptions in HTML files consistently began with the score information in the range of 100,200..,700, which we utilized as a criterion for selecting the data. To ensure a diverse selection of problems, we categorize the problems seven groups based on the score information. From the first five groups, we sample 60 problems per group. Since there are only 17 problems with scores of 600 and 700 in total, we include all of them, resulting in a total of 317 distinct problems.
For each problem, CodeNet provides multiple submissions, allowing us to capture a range of coding styles. Each submission is assigned a status, which can be: (i) accepted, (ii) compile-time error, (iii) runtime error, (iv) wrong answer, or (v) time limit exceeded. For each sampled problem, we select five submissions for each of the accepted, runtime error, and wrong answer statuses, resulting in a total of 15 (5x3) submissions per problem. This approach enables us to include a variety of coding styles and solutions, including incorrect ones. Note that there were no submissions with compile-time errors in our sample as Python is an interpreted language, not compiled. In addition, we exclude submissions with a time limit exceeded status, as we could not consistently find five such submissions for each problem. Moreover, since we cannot ascertain if these solutions are inherently incorrect or simply inefficient, we leave the inclusion of slow coding solutions for future work.
Overall, we sampled a total of 4,755 (317 x 5 x 3) code snippets from the CodeNet dataset, covering various coding problems and a wide range of correct and incorrect solutions written by different individuals.
3.2 Creating AI-Generated Code Dataset
To build the AI-generated code dataset, we employ three large language models (LLMs): i) CodeLlama Roziere et al. (2023) (34B parameters), ii) Codestral team (2024) (22B parameters), and iii) Gemini Team et al. (2024) (1.5 Flash). Using the human-authored code samples, we generate AI code enabling us to examine the distinctions between human-authored and AI-generated code. Specifically, for each coding problem and LLM, we generate three variations: i) generating code from scratch for a given problem, ii) fixing human-written code that results in a runtime error, and iii) correcting human-written code with incorrect output. This approach allows us to cover a range of AI-driven code-generation scenarios.
@@ 3.2 Creating AI-Generated Code Dataset
Across all trials,we generated 2,852 (317x3x3) code snippets, covering three distinct LLMs, three usage scenarios, and 317 coding problems.
3.3 Post-processing
@@
3.4 Brief Statistics
In order to provide more insight about our dataset, we provide sample code for the counting trailing zero problem. Specifically, the problem is as follows.
@@ 3.4 Brief Statistics
To examine the differences between AI-generated and human-written code, we calculated the average number of lines, blank lines, comments, and function definitions for human-written codes and for the outputs of the three LLMs separately. For AI-generated samples, we focused solely on codes generated from scratch to better capture distinctions in LLM-generated code. The results are presented in Figure 1.
@@
We observe that human-written codes tend to be longer compared to those generated by LLMs. Among the LLMs, CodeLlama defines more functions than both humans and the other models. Codestral, on the other hand, uses blank lines and comments more frequently, aligning more closely with human-written code in this regard.
@@ 4.1.1 Experimental Results
We assess the performance of the baseline models in our test test which includes all three LLMs and covers three different LLM usage scenarios. Table 3 shows the results.
@@
Lastl, we explore how the models’ performance is affected by the code generation scenario. In particular, we divide the test based on the generation scenario and calculated the recall for each scenario. The results are shown is Table 5.
@@ 5 Conclusion
In this work, we created AIGCodeSet, a dataset for the AI-generated code detection problem, focusing on Codestral, Codellama, and Gemini Flash 1.5 which are not well-studied in the literature. From the CodeNet dataset, we selected 317 problems and included five human submissions for each of the code status: Accepted, Runtime Error, and Wrong Answer, yielding 4,755 human-written code samples in total. Next, we utilized the mentioned LLMs in three approaches: i) we used them to generate code based solely on the problem description filei ii) we employed them to fix human-written code with runtime errors using the problem description file, iii) we prompted them to correct human-written code that resulted in wrong answers, again with the problem description file. After filtering out improper outputs, we obtained a total of 2,828 AI generated code samples. Eventually, AIGCodeSet consists of 7,583 (=4,755 + 2,828) data code samples. Furthermore, we conduct experiments on our dataset with baseline AI-generated code detection systems and share their results for different LLMs and LLM usage scenarios.
In the future, we plan to extend our dataset covering more programming languages, LLMs, and programming tasks to make our dataset more useful for effective training and reliable evaluation. We also plan to cover more usage scenarios such as blended codes where LLMs are used to generate a portion of the code. Furthermore, we will conduct a user study across students and software developers on how they use LLMs to generate code to identify realistic LLM usage scenarios.

### v2

@@ Abstract
While large language models provide significant convenience for software development, they can lead to ethical issues in job interviews and student assignments. Therefore, determining whether a piece of code is written by a human or generated by an artificial intelligence (AI) model is a critical issue. In this study, we present AIGCodeSet, which consists of 2.828 AI-generated and 4.755 human-written Python codes, created using CodeLlama 34B, Codestral 22B, and Gemini 1.5 Flash. In addition, we share the results of our experiments conducted with baseline detection methods. Our experiments show that a Bayesian classifier outperforms the other models.
@@ 1 Introduction
Artificial Intelligence (AI)-powered large language models are playing an important role in software development, significantly enhancing productivity. Beyond general-purpose chat-based models, specialized models such as CodeLlama, Codestral, and StarCoder have been specifically trained for code generation, enabling more efficient and automated software engineering workflows.
However, the widespread adoption of AI-generated code also introduces critical challenges, particularly concerning reliability and trustworthiness Wang et al. (2024); Rabbi et al. (2024); Ambati et al. (2024). These codes can sometimes contain unoptimized algorithms, inefficient structures, or faulty logic.
On the other hand, ethical violations such as candidates trying to deceive employers by using AI tools in job interviews [^1] https://fortune.com/2024/10/23/elon-musks-xai-cofounder-calls-out-cheating-interviewee-and-now-employers-are-outing-the-ai-tools-being-abused-by-savvy-job-seekers/ and students having AI models do their homework are also a major concernDeng and Joshi (2024).
For all these reasons, distinguishing AI-generated codes from human-written codes has become a critical requirement for software security and ethical standards.
In this study, we create a dataset to analyze AI-generated codes and compare them with human-written codes. Within the scope of our study, we compiled human-written codes for selected problems from the IBM CodeNet dataset and generated Python codes using three different models (CodeLlama, Codestral, and Gemini) in three different AI model usage scenarios in response to the same problems.
We perform various preprocessing steps on the obtained dataset to ensure the correctness and consistency of the codes. Then, we analyze structural features such as code length, comment lines, and function definitions to show whether the AI-generated codes exhibit certain patterns compared to human-written codes. Finally, we measure the performance of various basic methods such as Ada vectors, TF-IDF vectors, and Bayesian classifier to distinguish between human-written and model-generated codes.
Our experimental results show that the Bayesian classifier provides the highest sensitivity overall. However, we observe that the detection success of the models varies significantly depending on the major language model and usage scenario used.
The main contributions of our work can be summarized as follows:
  1. 1.
    We have created a comprehensive dataset containing a total of 4,755 human-written and 2,828 AI-generated codes that can be used in the detection of AI-generated codes. We share our dataset and codes to support research in this field and to reproduce our experimental results.  [^2] Data and code links are hidden due to confidential peer review.
  2. 2.
    We measured the baseline performances of various machine learning models on our dataset.
@@ 2 Related Work
Studies on AI-generated codes vary in terms of the target programming language, the purpose of creating the dataset, the size of the dataset, and the models selected and the source from which the human data is collected. For example, in terms of programming language, there are studies on various languages such as Python Idialu et al. (2024); Oedingen et al. (2024); Pan et al. (2024), Java Suh et al. (2024); Pham et al. (2024), Ruby Xu and Sheng (2025), Go Gurioli et al. (2024), and PHP Wang et al. (2023).
Datasets for detecting AI-generated codes also require human-written codes. Similar to our work, studies have generally found problem sets defined from different sources and a set of human-written codes that are solutions to these problems, and then had various LLMs generate codes for this problem. Studies in the literature have obtained human-written codes from various sources such as CodeChef Idialu et al. (2024), Kaggle, Quescol Pan et al. (2024), CodeNet Xu and Sheng (2025, 2024) and Rosetta Code project Gurioli et al. (2024). In our study, we also obtained human-written codes from CodeNet.
Moreover, GPT-based models such as ChatGPT Pan et al. (2024); Suh et al. (2024) and GPT-4 Suh et al. (2024); Pham et al. (2024); Xu and Sheng (2025) are widely utilized in the literature. Additionally, models like StarCoder Idialu et al. (2024); Gurioli et al. (2024), CodeBision Pham et al. (2024), and Gemini Suh et al. (2024) have also been employed in various studies. In our research, we utilize the CodeLlama, Codestral, and Gemini models.
Studies in the literature Pan et al. (2024); Oedingen et al. (2024) have focused on detecting code generated from a given problem. However, in this study, in addition to this problem, we also considered the cases of detecting runtime errors or codes that produce incorrect outputs, which are fixed using AI.
@@
3 Our Dataset
In this section, we share the process of collecting codes written by humans, the stages of code generation by large language models, the operations applied to the obtained data, and the statistics regarding the dataset.
3.1 Collecting Human Written Codes
The first step in creating the dataset is to identify coding problems. In this study, we used the CodeNet dataset [^3] https://developer.ibm.com/exchanges/data/all/project-codenet/ prepared by IBM and previously used in other studiesXu and Sheng (2025, 2024). CodeNet is a dataset that includes 4,000 different coding problems and solutions written by humans in different programming languages such as Python, C++, Java, Ruby. Sometimes, human-written codes in CodeNet pass all tests successfully, while sometimes they result in compile-time errors, runtime errors, time limit exceeds, or incorrect output errors.
In this study, we used the problems in CodeNet’s 800-problem subset called "Python Benchmark". This subset includes 800 problems that only contain human-written codes that have passed all tests successfully. While examining the problem descriptions, we noticed that most of them were assigned a score (in description text). Due to time and budget constraints, we took a sample of these problems, taking into account the scores in the problem descriptions. Specifically, to ensure as even a distribution as possible between the scores, we first sampled 60 problems in each of the first five groups (score ranges 100, 200, 300, 400, 500), making a total of 300 problems (5x60). Since there were only 17 problems with higher scores (600 and 700), we included these problems directly in the dataset, resulting in a final number of 317 problems.
All human-written code in CodeNet has labels indicating its status: i) Accepted, ii) Runtime Error, iii) Wrong Answer, iv) Compile-time Error, and v) Time-Limit Exceed error. To increase the diversity in our dataset, we collected human-written codes with different labels for each coding problem we pulled from the "Python Benchmark" section using CodeNet’s "Full (Original) Dataset" and "Full Dataset, Metadata Only" dataset. Specifically, we collected five code samples from each of the codes with accepted labels, runtime errors, and wrong outputs, creating a dataset of 4,755 (317 x 5 x 3 = 4,755) human-written codes in total. Since Python is an interpreted language, there is no code with the compile error tag. Furthermore, since the code with timeout errors is rare for the selected problem set, we did not include the code with these errors in our dataset.
3.2 Generating Code Snippets
@@
3.3 Filtering
@@
3.4 Features & Statistics
To offer a deeper understanding of our dataset, we provide sample code for the counting trailing zero problem. Specifically, the problem is as follows.
@@ 3.4 Features & Statistics
To analyze the distinctions between AI-generated and human-written code, we calculated the average number of lines, blank lines, comments, and function definitions for human-written codes and for the outputs of the three LLMs separately. For AI-generated samples, we focused solely on codes generated from scratch to better capture distinctions in LLM-generated code. The results are presented in Figure 1.
@@
Human-written code is typically more extensive than code generated by LLMs. Among the LLMs, CodeLlama defines more functions than both humans and the other models. Codestral, on the other hand, uses blank lines and comments more frequently, aligning more closely with human-written code in this regard.
@@ 4.1.1 Experimental Results
We assess the performance of the baseline models in our test set which includes all three LLMs and covers three different LLM usage scenarios. Table 3 shows the results.
@@
Lastly, we explore how the models’ performance is affected by the code generation scenario. In particular, we divide the test based on the generation scenario and calculated the recall for each scenario. The results are shown is Table 5.
@@ 5 Conclusion
In this study, we developed AIGCodeSet, a dataset for the AI-generated code detection problem, focusing on Codestral, Codellama, and Gemini Flash 1.5 which are not well-studied in the literature. From the CodeNet dataset, we selected 317 problems and included five human submissions for each of the code status: Accepted, Runtime Error, and Wrong Answer, yielding 4,755 human-written code samples in total. Next, we utilized the mentioned LLMs in three approaches: i) we used them to generate code based solely on the problem description filei ii) we employed them to fix human-written code with runtime errors using the problem description file, iii) we prompted them to correct human-written code that resulted in wrong answers, again with the problem description file. After filtering out improper outputs, we obtained a total of 2,828 AI generated code samples. Eventually, AIGCodeSet consists of 7,583 (=4,755 + 2,828) data code samples. Additionally, we perform experiments on our dataset using baseline AI-generated code detection systems and present their results across various LLMs and usage scenarios.